OpenStudy (anonymous):

**WILL MEDAL** Which ordered pair is the solution to the system of equations? -3x+4y=-20 x+4y=-4 A. (8, 1) B. (−12, 2) C. (−4, −8) D. (4, −2)

OpenStudy (fibonaccichick666):

so two ways, one just plug in the points and see if they work, 2 solve the system. Would you like to learn how?

OpenStudy (anonymous):

Um duh

OpenStudy (anonymous):

Lol jk

OpenStudy (fibonaccichick666):

ok, well, I have to check sometimes you get just answer people and I'm sick of wasting my time on them

OpenStudy (fibonaccichick666):

So, ready to work with me?

OpenStudy (anonymous):

Yea ig

OpenStudy (fibonaccichick666):

ok, so, you want to pick any equation, and decide whether you want to solve for y or x

OpenStudy (fibonaccichick666):

which do you prefer

OpenStudy (anonymous):

y first

OpenStudy (fibonaccichick666):

ok, so I recommend solving the top equation for y first then, do you know how to do that?

OpenStudy (anonymous):

Not really

OpenStudy (fibonaccichick666):

ok, np, so we are going to solve \[-3x+4y=-20\] for y. Our first step is to "undo" the addition of \(-3x\) on the left side do you have any ideas on what we can do?

OpenStudy (anonymous):

Add 3 to both sides?

OpenStudy (fibonaccichick666):

well, almost, we need to get rid of the -3x, so, we add +3x to both sides, can you try that for me please?

OpenStudy (anonymous):

4y=-23?

OpenStudy (fibonaccichick666):

well, we can't add -20 and 3x because they are not like terms, so we literally just say \[4y=-20+3x\] Does that make sense?

OpenStudy (anonymous):

Oh yes

OpenStudy (fibonaccichick666):

ok, so now, what's your next step to get y by itself?

OpenStudy (anonymous):

So 4y=-23x?

OpenStudy (fibonaccichick666):

nope, since -20 doesn't have an x, we cannot add an x term to it.(Just an fyi, if we did -20+3=-17)

OpenStudy (fibonaccichick666):

You are using common core eh? I'll be back in like 2 minutes, I have to go take out a load of laundry, but I'll teach you as soon as I get back. Can you please watch this video on solving a system of equations for me while I'm gone? https://www.khanacademy.org/math/algebra/systems-of-eq-and-ineq/fast-systems-of-equations/v/solving-linear-systems-by-substitution

OpenStudy (anonymous):

So would it be C?

OpenStudy (fibonaccichick666):

I don't think that -4+4*-8=-4 Do you?

OpenStudy (fibonaccichick666):

So, did you check out the video?

OpenStudy (anonymous):

Yea now I think it's b

OpenStudy (fibonaccichick666):

does -3*-12+4*2=-20?

OpenStudy (anonymous):

No it's 44

OpenStudy (fibonaccichick666):

so that can't be the answer since it doesn't wok for both equations, since we are currently doing a guess and check style, try A, Does it work?

OpenStudy (anonymous):

Yes?

OpenStudy (fibonaccichick666):

try it and show me

OpenStudy (fibonaccichick666):

put it into the second and see if it works

OpenStudy (fibonaccichick666):

does 8+4*1=-4?

OpenStudy (anonymous):

No 12

OpenStudy (fibonaccichick666):

ok, so does A work?

OpenStudy (anonymous):

No

OpenStudy (fibonaccichick666):

ok, so try our last one, Does D work, test it in both equations and show me here

OpenStudy (anonymous):

Finally answer-->D

OpenStudy (fibonaccichick666):

Does it work? Show me, we haven't solved anything

OpenStudy (anonymous):

I gtg now

OpenStudy (anonymous):

But thx anyways

OpenStudy (fibonaccichick666):

....

OpenStudy (fibonaccichick666):

yw I guess